find the area of a parallelogram whose base is 6 1/4 inches and whose altitude is 7 2/3 inches
Answer by stanbon(48558)

Area = (base)(altitude)
Area = (25/4)(23/3)
Area = 575/12= 41.916....
Cheers,
stan H